### v2

# Active-Absorbing Phase Transitions in the Parallel Minority Game

Abstract
The Parallel Minority Game (PMG) is a synchronous adaptive multi-agent model that exhibits active-absorbing transitions characteristic of non-equilibrium statistical systems. We perform a comprehensive numerical study of the PMG under two families of microscopic decision rules: (i) agents update their choices based on instantaneous population in their alternative choices, and (ii) threshold-based activation that activates agents movement only after overcrowding density crossing a threshold. We measure time-dependent and steady state limits of activity $A(t)$, overcrowding fraction $F(t)$ as functions of the control parameter $g=N/D$, where $N$ is the number of agents and $D$ is the total number of sites. Instantaneous rules display mean-field directed-percolation (MF-DP) scaling with $\beta\approx 1.00$, $\delta\approx 0.5$, and $\nu_{\parallel}\approx 2.0$. Threshold rules, however, produce a distinct non-mean-field universality class with $\beta\approx 0.75$ and a systematic failure of dynamical scaling. We show that thresholding acts as a relevant perturbation to the critical behavior of the model. The results highlight how minimal cognitive features at the agent level fundamentally alter large-scale critical behavior in socio-economic and active systems.
preprint: AIP/123-QED
Collective behavior in adaptive multi-agent systems often displays emergent macroscopic phenomena that are qualitatively distinct from the microscopic decision rules of individual agents, a theme central to both statistical physics and socio-economic modelingChakrabarti et al. (2006, 2023); Sen and Chakrabarti (2014). A paradigmatic example is the Minority Game (MG), introduced as a minimal model of competitive decision-making under bounded rationality Arthur (1994); Challet and Zhang (1997, 1998); Challet and Marsili (1999), and subsequently developed into a general framework for studying coordination, resource allocation, and market-like dynamics Challet et al. (2005); Chakraborti et al. (2015, 2011). Extensions of the MG, including spatial and local variants, have revealed rich collective behavior arising from interaction topology and information structure Challet et al. (2001); Moelbert and De Los Rios (2002); Huang et al. (2012). Minority-game-inspired models have been applied to wealth inequality, herding control, and social efficiency in competitive environments Ho et al. (2004); Zhang et al. (2016); Bottazzi and Devetag (2007); Dhar et al. (2011); Chakrabarti et al. (2009); Ghosh et al. (2010); Biswas et al. (2012). Synchronously updated systems that admit absorbing states (dynamics ceases entirely), known as active-absorbing transitions, are a cornerstone of non-equilibrium statistical mechanics, with directed percolation (DP) emerging as the generic universality class under broad conditions Henkel et al. (2000); Ódor (2004); Lübeck (2004); Henkel et al. (2008). Recently, the Parallel Minority Game (PMG) has been introduced as a multi-choice, globally coupled extension of the MG, in which agents switch synchronously between two fixed options drawn from a larger set Biswas and Mandal (2021); Vemula and Biswas (2025). The resulting non-conserved population dynamics and shared resource constraints naturally generate absorbing states and critical behavior. Understanding how adaptation mechanisms influences the system efficiency becomes very importantSysi-Aho et al. (2004). Here, the contrast between instantaneous updates and threshold-based activation not only determines the universality class of the transition but also offers a controlled framework to probe how minimal cognitive constraints function as relevant perturbations in non-equilibrium socio-economic systems.
I Introduction
Competitive multi-agent systems often exhibit emergent macroscopic behavior that is qualitatively different from individual decision rules. The Minority Game (MG) framework, introduced by Challet and Zhang  Challet and Zhang (1997), provides a minimal model of such competition: agents make binary choices and the minority side wins. The original MG and its variants, including the local MG where agents interact with a subset of the population, have been extensively studied for their phase transitions and collective behaviors Challet et al. (2001). In particular, the Local Minority Game introduced by Moelbert and De Los Rios considers agents placed on a lattice interacting with their nearest neighbors, leading to spatial correlations and an annealed disorder that differs from the quenched disorder of the original MG Moelbert and De Los Rios (2002). MG variants have been used to model financial markets, resource allocation, and collective dynamics in complex adaptive systems Challet and Marsili (1999); Chakraborti et al. (2015). The Parallel Minority Game (PMG), as detailed in recent extensions Biswas and Mandal (2021); Vemula and Biswas (2025); Xue et al. (2025), generalizes the MG to scenarios with $D>2$ sites, where each agent is restricted to switching between only two fixed choices among the $D$ available sites. The objective for the players remains the same viz., to be in the minority among the two available choices to them. But with $D>2$, each site can now contain agents having their alternate choice distributed among the various other sites. This overlap in choices creates coupled, parallel instances of the MG, introducing non-conserved population dynamics and heightened competition due to shared resources across agent groups.
The synchronous (parallel) update in PMG introduces non-equilibrium dynamics, including absorbing states where the system freezes and activity ceases. Absorbing-state transitions are central to non-equilibrium statistical mechanics; showing robust universality under broad conditions Henkel et al. (2008, 2000); Ódor (2004). (see Henkel et al. Henkel et al. (2008) and Hinrichsen Henkel et al. (2000)). In PMG, the frozen state, where no agents switch due to balanced populations, serves as the absorbing state, with global interactions approximating mean-field behavior.
In this study, we explore how microscopic decision rules shape the universality class of the PMG. In particular, we compare instantaneous decision rules with threshold-based rules. Our empirical finding is that instantaneous rules map to the expected mean field behavior, while threshold rules shift the system to a distinct universality class. This is consistent with earlier work on adaptive minority games, where threshold-based decision mechanisms were found to significantly alter system behavior and drive it toward different regimes Sysi-Aho et al. (2004).
II Model, Strategies and Observables
We consider $N$ agents distributed randomly and uniformly among $D$ sites and a second alternative site is also assigned to each agent randomly. Each agent is restricted to move between exactly the two fixed choices for them (e.g., agent $i$ toggles between sites $x_{i}$ and $y_{i}$), and every site is ensured to be at least of the choices of at least one of the agents i.e., there is no site that is not part of the game at all. The control parameter is
$g=\frac{N}{D},$
representing the average population density of the choices. All simulations are performed for $D=500$, and the value of $N$ is fixed when a particular $g$ value is chosen. The point of interest is the behavior of the system near $g=g_{c}=1$, where the average occupancy per site reaches unity, marking the onset of overcrowding and the active-absorbing transition. Below this point, on average, no agent will be in the majority location. The random allocation of the agents’ choices create overlaps in the choices (any given choice/site has agents whose alternate choice can be from any one of the rest D1D-1 choices). This situations create coupled minority sub-games. Each agent $i$ at time $t$ takes an action $a_{i}(t)\in\{0,1\}$, representing their current site choice: $a_{i}(t)=0$ implying no change in choice for the next step and $a_{i}(t)=1$ implying a switch to the alternate choice in the next step.
This model has been studied before Biswas and Mandal (2021) in the context of movement of infected agents when an infectious disease is spreading in different locations, and also subsequently in the context of finding efficient strategy for the agents to be in the minority in the limit of high population density (ggVemula and Biswas (2025). Of course, for high $g$ values, some agents will always be in the overcrowded sites, but a memory-based efficient strategy could maximize (within the strategies explored) the fraction of agents in the minority.
In this work, however, we look at the low density phase ($g$ close to one), where it is possible for $g\leq 1$ to reach a state where no agent is in overcrowded sites. But for $g>1$, the overcrowding can be represented by an active-absorbing transition. For characterization of such a transition, we then define the observables activity and overcrowding fraction, which quantify overcrowding and persistent competition in the model:
{align}
A(t)=1D∑_s=1^D max(n_s(t)-1,0),
F(t)=1D∑_s=1^D 1 {n_s(t)>1},
where $n_{s}(t)$ is the occupancy at site $s$, and $\mathbbm{1}\{\cdot\}$ is the indicator function.
II.1 Stochastic Strategy Formulation
For comparative analysis, we consider two stochastic strategies adapted from recent work on efficient PMG dynamics Vemula and Biswas (2025):
Strategy A: Instantaneous minority comparison
Agents use current population information from both choices and switches with a probability given below, only if $n_{x_{i}(t)}>n_{y_{i}(t)}$ and $n_{x_{i}(t)}>1$:
$p_{i}^{A}(t)=\frac{n_{x_{i}(t)}-n_{y_{i}(t)}}{2n_{x_{i}(t)}}$
where $y_{i}(t)$ is the alternate location. $p_{i}^{A}(t)=0$, if $n_{x_{i}(t)}\leq n_{y_{i}(t)}$ or $n_{x_{i}(t)}\leq 1$. This strategy directly implements minority seeking with current information, similar to memory-less updates.
Strategy B: Threshold comparison
Agents switch with a probability only if there is a positive deviation from the global average occupancy ($n_{x_{i}}(t)>g$ and $n_{x_{i}}(t)>1$):
$p_{i}^{B}(t)=\frac{n_{x_{i}}(t)-g}{2n_{x_{i}}(t)},$
where $x_{i}(t)$ is the current location of agent $i$, $n_{x_{i}}(t)$ is the population at that location, and $g=N/D$ is the global average occupancy. $p_{i}^{B}(t)=0$, if $n_{x_{i}}(t)\leq g$ or $n_{x_{i}}(t)\leq 1$. This strategy drives population toward uniform distribution but doesn’t consider minority status while making the decision on switching.
III Scaling Theory and Static Critical Behavior
The objective of the players in the Parallel Minority Game, as mentioned before, is to be in the minority between their two choices. The maximum utilization, therefore, is achieved when the fluctuation in the population among the different choices is low (same measure as the classical Minority Game problem, where there are only two choices). This assumes, of course, that all choices are nominally similar in terms of the number of player allocated, which is true since the players are assigned their choices randomly and uniformly (held fixed throughout the game). Now, when the number of players equal the number of choices (i.e., $g=1$), a situation would arise where each choice is occupied by one agent, which is an absorbing phase, meaning either strategy would not give a non-zero probability for any agent to switch. The dynamics becomes interesting, however, when $g>1$. Due to the strategies, the actual number of players who would now be in overcrowded locations (having population greater than one), can now be more than what is the minimum required (gNDgN-D). We quantify the scaling behavior of this population dynamics, using the definitions of activity and overcrowding fractions defined above, in the steady state limit, to explore the critical behavior.
We probe the steady-state values $A_{\infty}(g)$ and $F_{\infty}(g)$ for $g>g_{c}$ and extract the order-parameter exponent $\beta$ via
$A_{\infty}\sim(g_{c}-g)^{\beta},\qquad F_{\infty}\sim(g_{c}-g)^{\beta}.$
The finite-size scaling analysis requires careful treatment of system sizes and fitting ranges to avoid crossover effects near $g_{c}$. We employ ensemble averaging over multiple realizations to ensure statistical reliability.
Figure 1: Static scaling fits for the steady-state activity and overcrowding fraction near $g_{c}=1$. Top row: Instantaneous rules (strategy A) yield $\beta\approx 0.95\pm 0.05$, consistent with mean-field behavior . Bottom row: Threshold rules (strategy B) yield $\beta\approx 0.75\pm 0.05$, indicating a different universality class.
The data for strategy A display near-linear behavior on a log-log plot (see Fig. 1), with $\beta_{A}^{\rm inst}\simeq 0.95\pm 0.05$ and $\beta_{F}^{\rm inst}\simeq 0.96\pm 0.05$, confirming the expected mean field scaling  Henkel et al. (2008). However, in the strategy B, the data exhibit systematically lower slopes corresponding to $\beta\approx 0.75$ (consistently measured across both observables and different threshold variants), indicating a clear deviation from the expected mean field universality. This aligns with earlier findings that adaptation mechanisms in minority games can drive systems toward different behavioral regimes Sysi-Aho et al. (2004), where threshold-based strategies alter the fundamental dynamics.
IV Dynamical Scaling Framework
Dynamical scaling provides a more stringent test of universality than static exponents alone. For systems exhibiting absorbing phase transitions, the time-dependent order parameter follows the scaling form:
$A(t)=t^{-\delta}\,\mathcal{A}\!\big(t|g-g_{c}|^{\nu_{\parallel}}\big),$
where $\delta$ is the temporal decay exponent and $\nu_{\parallel}$ is the temporal correlation length exponent Ódor (2004). Similarly, for the overcrowding fraction,
$F(t)=t^{-\delta}\,\mathcal{F}\!\big(t|g-g_{c}|^{\nu_{\parallel}}\big),$
with the same exponents $\delta$ and $\nu_{\parallel}$, and scaling function $\mathcal{A}(x)$ sharing asymptotic behaviors with $\mathcal{F}(x)$. For the mean field limit, the expected exponents are $\delta_{\rm MF}=1$ and $\nu_{\parallel,{\rm MF}}=1$.
The scaling function $\mathcal{F}(x)$ has the asymptotic behavior: $\mathcal{F}(x)\sim\text{constant}$ for $x\ll 1$ (critical quenching), and $\mathcal{F}(x)\sim x^{\beta}$ for $x\gg 1$, ensuring consistency with static scaling $A_{\infty}\sim(g_{c}-g)^{\beta}$ Henkel et al. (2000).
We test this dynamical scaling for strategy A, which should conform to the mean field predictions. The data collapse procedure involves plotting $A(t)t^{\delta}$ against $t|g-g_{c}|^{\nu_{\parallel}}$ and assessing the quality of superposition across different $g$ values, using rainbow-colormapped curves for $g\in[0.75,1.15]$ (excluding $g_{c}$) (see Fig. 2).
Figure 2: Dynamical scaling collapses for instantaneous strategy (strategy A) using mean field exponents $(\delta,\nu_{\parallel})=(0.5,2.0)$. Data Collapse shows superposition into master curve across $g$ values. This confirms single-parameter finite-time scaling and mean field universality Vemula and Biswas (2025).
The data collapses for strategy A indicate consistency with the predicted exponents $\delta=0.5$ and $\nu_{\parallel}=2.0$. This confirms that the critical behavior corresponding to strategy A belong to the mean field universality class, as expected from theoretical considerations Henkel et al. (2008).
Threshold-based decision rules (strategy B) introduce fundamental differences from instantaneous rules. Unlike instantaneous rules where decisions depend only on current population differences, threshold rules require agents to look at the average population (ignoring the majority/minority status) before switching states.
Figure 3: Dynamical scaling collapses for threshold strategies (strategy B) using non-mean field exponents. (a-b) Supercritical regime: $(\delta,\nu_{\parallel})=(1.0,0.75)$. (c-d) Subcritical regime: $(\delta,\nu_{\parallel})=(0.2,1.0)$. Log-log plots reveal systematic fanning when using mean field exponents, with partial up/down overlaps indicating multiple timescales. Failure across variants and observables confirms deviation from mean field universality, suggesting a distinct universality class.
The deviations (from strategy A and mean field expectations) are apparent in the attempts to data collapse (see Fig. 3). Attempts to collapse the data using mean field exponents produce systematic curvature and poor superposition. We find different scaling regimes: the supercritical region requires exponents $\delta=1.0$ and $\nu_{\parallel}=0.75$, while the subcritical region requires $\delta=0.2$ and $\nu_{\parallel}=1.0$. This persistent failure, combined with the consistently different static exponent $\beta\approx 0.75$, provides compelling evidence for a distinct universality class for the model following strategy B.
V Discussions and conclusions
We have presented a comprehensive numerical study of phase transitions in the Parallel Minority Game under two distinct decision paradigms. Our key findings are that instantaneous decision rules (strategy A) yield an active-absorbing transition in terms of the fraction of agents not in the minority or the fraction of sites where over-crowding happens, with critical exponents ($\beta\approx 0.95$, $\delta\approx 0.5$, $\nu_{\parallel}\approx 2.0$) (see Fig. 1 and  2) and consistent with mean field values for such exponents. But the threshold-based rules (strategy B) produce a distinct universality class characterized by $\beta\approx 0.75$ (see Fig. 1) and systematic failure of dynamical scaling, with evidence of multi-timescale relaxation and different exponent values in supercritical ($\delta=1.0$, $\nu_{\parallel}=0.75$) (see Fig. 3) (a-b) and subcritical ($\delta=0.2$, $\nu_{\parallel}=1.0$) (see Fig. 3) (c-d) regimes. The deviation stems from the fundamental nature of thresholding, which acts as a relevant perturbation in the renormalization group sense.
We should also note the possible link of the model dynamics to a frustrated spin-glass like system Sherrington and Kirkpatrick (1975). In such systems, due to both ferromagnetic and anti-ferromagnetic coupling between (Ising) spins, the ground state becomes highly degenerate and difficult to find, the energy landscape being highly corrugated. The system dynamics can often lead to a local minimum rather than the global one. Similarly, in the present system, decrease in population in one site necessarily increases the same in other sites, making the ’ground state’ or optimized distribution of crowds (presumably equal population in each site) a difficult goal. The effect of such ’frustration’ in the critical exponent values studied above, and the effectiveness of heuristic techniques (see e.g., Das et al. (2025)) to find the ground states here could be an important direction.
In conclusion, our results demonstrate that minimal cognitive features such as decision thresholds can fundamentally alter universality classes in adaptive multi-agent systems. This has several important implications.
Real-world decision makers often exhibit inertia and threshold behavior-consumers wait for significant price changes before switching brands, investors require substantial evidence before changing portfolios, and firms delay responses until profit differentials exceed certain thresholds. Our findings suggest that such micro-level cognitive features can dramatically affect macroscopic collective behavior and phase transition characteristics.
Code Availability
The simulation code for this work is available at https://github.com/AryanTyagi-Complexity/Parallel-Minority-Game